[Brahms-l] Shift report 20050618 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Jun 18 2005 - 08:05:04 EDT
Shifter leader was AG
Shifters were 


1. FS moved from 4 to 2.3 deg after run 15601.
Since ramp of beam the background levels have been high.

2. Fill pattern: 95x95_P4,Polarizations: B/Y: -47.47+/-2.43,-46.79+/-2.25, with FS on 4 deg:<br>
Run #15596-15600<br>
<br>
Moved FS to 2.3 deg.<br>
Pedestal run #15602, Non pedestal run #15603, Pedestal run #15604.<br>
Pedestal run #15602 was recorded in daq and tpcdped as having FS at 4 deg. This was the reason that there was made an other pedestal run (#15604), which has corrected FS to 4 deg.<br>
<br>
Fill pattern: 106x106_P1,Polarizations: B/Y: -41.80+/-2.37,-37.87+/-2.30, with FS on 2.3 deg:<br>
Run #15605-15608<br>

Due to printer out of toner the supermon data (run #15599,15605-15608) was stores in:<br>
/opus/home/operator/public_html/MonitorSummary/run05/<br>

Run 15601 should be discarded because it is only ~10 min long.<br>

3. Still voltage violation on T3, A3-xu: demand 0, readout 21.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jun 18 08:05:16 2005

This archive was generated by hypermail 2.1.8 : Sat Jun 18 2005 - 08:05:27 EDT